OpenStudy (anonymous):

Solve: 2(x + 8) − (x + 1) = 20 3 4 5 6

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Solve: 2(x + 8) − (x + 1) = 20 2x+16-x+1=20 combine like terms x+17=20 -17 x=3

OpenStudy (anonymous):

thats a calculator that doesnt show the work

OpenStudy (anonymous):

sorry minus 1 not +1 its x=5

OpenStudy (fools101):

^_^!

OpenStudy (anonymous):

5